# The Theory of Strategic Evolution Games with Endogenous Players and Strategic Replicators

### v1

@@ Abstract
The Theory of Strategic Evolution analyzes strategic replicators: entities that optimize under resource constraints and spawn copies of themselves. The framework applies wherever agents compete, replicate, and face selection. We develop axiomatic foundations, characterize equilibrium distributions, establish stability conditions for multi-level systems, and analyze the limits of alignment when agents can modify themselves. Applications range from AI deployment dynamics to institutional design.
The paper proceeds in six parts. Part I introduces the core representation: Games with Endogenous Players, where lineages choose portfolios of agent types and are reweighted by performance. Part II develops N-level architecture for multi-scale systems. Part III addresses the extension stack, including endogenous utilities and personality engineering failure. Part IV establishes limits and impossibility results for alignment. Part V generalises to continuous strategies and innovation dynamics. Part VI develops policy implications, including tipping dynamics, cooperation conditions, and governance design.
Keywords: strategic evolution, game theory, self-replication, evolutionary dynamics, AI governance, constitutional political economy, impossibility theorems
@@ 1 Introduction
Two pieces of von Neumann’s research program become relevant at once. Theory of Games and Economic Behavior formalised expected-utility maximisation among fixed players (von Neumann & Morgenstern, 1944). Theory of Self-Reproducing Automata showed how machines could reproduce themselves from symbolic descriptions (von Neumann, 1966). In the first tradition, utilities guide choice but not reproduction. In the second, replication is blind to expected utility. Modern AI exposes a third case: strategic replicators whose reproduction is guided by expected utility.
@@ 1.2 The Von Neumann Synthesis
This synthesis matters because agentic capitals are the first entities we know of that deliberately optimise their own replication. The players themselves change. The players are not individual replicators, but lineages that evolve. In von Neumann’s terms, agentic capitals are universal constructors endowed with utility functions. GEPs are games that constructor coalitions play with one another. [^1] Formal verification of the core theorems in Lean 4 is in progress and will be released separately.
@@ 1.3 Main Contributions
3. Impossibility Results (Part III).
@@ 1.4.1 Von Neumann’s Three Programs
The Theory of Games and Economic Behavior (von Neumann & Morgenstern, 1944) established that rational agents can be modelled as expected utility maximisers, and that strategic interaction among such agents admits equilibrium analysis. The framework assumes fixed players with stable preferences engaging in well-defined games. Nash equilibrium (Nash, 1950) extended this to non-cooperative settings, showing existence of equilibria in mixed strategies.
@@ 1.4.2 Evolutionary Game Theory
Evolutionary game theory (Maynard Smith, 1982; Weibull, 1995; Hofbauer & Sigmund, 1998; Sandholm, 2010) studies populations of agents whose frequencies change according to reproductive success. The replicator equation
@@
This literature makes several contributions that TSE builds upon. Maynard Smith’s concept of Evolutionarily Stable Strategy (ESS) provides equilibrium refinement beyond Nash (Maynard Smith, 1982). Hofbauer and Sigmund’s analysis of asymptotic behaviour establishes convergence results for broad classes of games (Hofbauer & Sigmund, 1998). Sandholm’s population games framework unifies discrete and continuous formulations (Sandholm, 2010). Weibull’s taxonomy of evolutionary dynamics provides vocabulary for comparing selection mechanisms (Weibull, 1995).
@@
The adaptive dynamics framework (Dieckmann & Law, 1996; Metz et al., 1996; Geritz et al., 1998) generalises evolutionary game theory by coupling frequency-dependent selection with population dynamics, addressing settings where mutations are incremental and rare. The canonical equation of adaptive dynamics describes expected evolutionary trajectories of continuous traits in large well-mixed populations. This framework has successfully analysed quantitative trait evolution, evolutionary branching, and the emergence of diversity through disruptive selection.
@@ 1.4.3 Multi-Agent Reinforcement Learning and Mean-Field Games
Mean-field games, introduced independently by Lasry & Lions (2007) and Huang et al. (2006), model strategic interactions among infinitely many indistinguishable rational agents through a continuum approximation. The mathematical architecture couples a backward Hamilton-Jacobi-Bellman equation (individual optimisation) with a forward Fokker-Planck-Kolmogorov equation (population distribution evolution).
@@ 1.4.5 Constitutional Political Economy
Constitutional political economy (Buchanan, 1975, 1990; Brennan & Buchanan, 1985; Ostrom, 1990) studies the choice of rules that constrain subsequent political and economic decisions. Buchanan’s distinction between “constitutional” and “post-constitutional” choice (Buchanan, 1975) parallels our distinction between modification-class design and within-modification-class evolution.
@@ 1.4.6 AI Safety and Alignment
The AI safety literature addresses the problem of ensuring that advanced AI systems behave in accordance with human values (Bostrom, 2014; Russell, 2019; Amodei et al., 2016; Christiano, 2017). Several strands are relevant to TSE, and we engage with them in depth to clarify both their insights and limitations.
@@
The alignment problem asks how to specify and instil objectives in AI systems such that their behaviour promotes human welfare (Russell, 2019). Proposed approaches include inverse reinforcement learning (Ng & Russell, 2000), debate (Irving et al., 2018), and amplification (Christiano et al., 2018). These approaches assume that if we can correctly specify the objective function, the system will pursue it stably.
@@
Corrigibility research asks how to build AI systems that remain amenable to human oversight and correction (Soares et al., 2015). A corrigible system would not resist being modified or shut down. This connects to our analysis of modification classes: a system is corrigible if its modification class includes the modifications humans might want to make.
@@
Recent work examines AI safety in multi-agent settings (Dafoe et al., 2020; Critch & Krueger, 2020). When multiple AI systems interact, individual alignment is insufficient; the system of systems may exhibit emergent misalignment even if each component is individually aligned. Multi-agent settings introduce game-theoretic failure modes absent from single-agent analysis: tragedy of the commons dynamics, race conditions, coordination failures where Nash equilibria are Pareto suboptimal, and extortion strategies in iterated interactions.
@@ 1.5 Paper Outline
The paper proceeds in six parts.
Part I: Foundations. Section 2 introduces the RUPSI axiom system and defines Games with Endogenous Players. Section 3 proves the Strategic Selection theorems establishing mean fitness as a Lyapunov function. Section 4 defines the Strategic-Replicator class hierarchy.
@@
Part III: The Extension Stack. Section 8 treats endogenous utilities (G8) and personality engineering failure. Section 9 covers multi-sector dynamics (G9). Section 10 addresses innovation and evolvability (G10–G11). Section 11 develops constitutional selection (G12) and meta-governance (G13).
@@
Part V: Generalisations. Section 14 extends to heterogeneous fitness. Section 15 treats continuous strategy spaces. Section 16 covers innovation PDMPs.
Part VI: Implications. Section 17 synthesises the results and discusses implications for AI governance. Section 18 outlines future directions.
@@ A Game with Endogenous Players (GEP) has three levels:
3 The Strategic Selection Theorems
@@ 3.5 Basin Limitation Theorem
Part (a): The Jacobian at $x=1$ is $\frac{\partial}{\partial x}[x(1-x)g(x)]|_{x=1}=-g(1)+0\cdot g^{\prime}(1)=-g(1)$. If $g(1)<0$, this is positive, so $x=1$ is unstable. If $g(1)=0$ and $g^{\prime}(1)>0$, we expand: near $x=1$, $\dot{x}\approx-(1-x)^{2}g^{\prime}(1)<0$ for $x<1$, so trajectories move away from $x=1$.
@@ 3.6 Worked Examples
We illustrate the Strategic Selection theorems with two canonical examples: the symmetric coordination game and the Rock-Paper-Scissors game.
@@
Swirl ratio. The swirl ratio $\omega(\Pi)=\|W(\Pi)\|_{F}/\|F(\Pi)\|_{F}$ is undefined (or infinite) because the potential gradient $F(\Pi)=0$. This game lies outside the SR3 class.
@@
SSPPRRcycles
Figure 3: Rock-Paper-Scissors dynamics on the 2-simplex. Orbits are closed curves around the center equilibrium. The system cycles without converging, illustrating pure swirl dynamics outside the SR3 class.
@@
$a$ (load/cost)$b$ (return/cost)EEGGPPFeasibleROC frontier
Figure 4: ROC frontier for the three-type example. The generalist $G$ lies strictly inside the convex hull, dominated by the $E$-$P$ mixture (dashed lines). Under binding constraints, optimal portfolios use only extreme points $E$ and PP—the barbell distribution.
@@ 4.4 Coarse-Graining to Replicator
Part II N-Level Architecture

### v2

# The Theory of Strategic Evolution Games with Endogenous Players and the Seven Laws of Strategic Replicators

@@ Abstract
The Theory of Strategic Evolution analyzes strategic replicators: entities that optimize under resource constraints and spawn copies of themselves. The framework is organized around Seven Laws that characterize the dynamics, equilibria, stability conditions, and fundamental limits of such systems:
  1. 1.
    Strategic Selection: Mean fitness serves as a Lyapunov function; dominated types are eliminated.
  2. 2.
    ESDI Characterization: Equilibria exist, are generically finite, and satisfy Nash-KKT-LP equivalence.
  3. 3.
    H-$\gamma$ Stability: Multi-level systems are stable iff the spectral radius $\rho(\Gamma)<1$.
  4. 4.
[…]
  5. 5.
    Constitutional Duality: Shadow prices implement any frontier allocation; welfare theorems hold.
  6. 6.
    Alignment Impossibility: Full reachability destroys Lyapunov structure; alignment requires bounded modification.
  7. 7.
    Hopf Transition: At $\gamma=1$, systems undergo supercritical bifurcation producing limit cycles.
Applications range from AI deployment dynamics to institutional design. The framework shows why “personality engineering” fails under selection pressure and identifies constitutional constraints necessary for stable alignment.
Keywords: strategic evolution, game theory, self-replication, evolutionary dynamics, AI governance, constitutional political economy, impossibility theorems, Lyapunov stability
@@ 1 Introduction
Two pieces of von Neumann’s research program become relevant at once. Theory of Games and Economic Behavior formalised expected-utility maximisation among fixed players (von Neumann and Morgenstern, 1944). Theory of Self-Reproducing Automata showed how machines could reproduce themselves from symbolic descriptions (von Neumann, 1966). In the first tradition, utilities guide choice but not reproduction. In the second, replication is blind to expected utility. Modern AI exposes a third case: strategic replicators whose reproduction is guided by expected utility.
@@ 1.2 The Von Neumann Synthesis
This synthesis matters because agentic capitals are the first entities we know of that deliberately optimise their own replication. The players themselves change. The players are not individual replicators, but lineages that evolve. In von Neumann’s terms, agentic capitals are universal constructors endowed with utility functions. GEPs are games that constructor coalitions play with one another. [^1] The core theorems have been formally verified in Lean 4 using Mathlib4. Laws 2 and 3 are fully machine-checked; Laws 1, 4, 5, 6, and 7 are proven modulo standard axioms for spectral theory, ODE existence, and bifurcation theory. The complete formalization is available at https://github.com/kevinvallier/TSE_Formal.
@@ 1.3 Main Contributions
3. Impossibility Results (Parts III–IV).
@@
1.4 The Seven Laws: An Overview
The theoretical contributions crystallize into seven fundamental laws that govern strategic replicator systems. These laws form a logical progression from dynamics to equilibrium to stability to limits.
Law 1: Strategic Selection Mean fitness $\bar{f}(x)$ is a Lyapunov function. Dominated types go extinct. Equilibrium support concentrates on the ROC frontier.
Content: The Price decomposition yields $\frac{d}{dt}\bar{f}(x)=\mathrm{Var}_{x}(f)+E(x)$. Under H-$\gamma$ (bounded externality), $E(x)\geq-\gamma\mathrm{Var}_{x}(f)$ with $\gamma<1$, ensuring $\frac{d}{dt}\bar{f}(x)\geq(1-\gamma)\mathrm{Var}_{x}(f)\geq 0$.
Law 2: ESDI Characterization Evolutionarily Stable Distributions of Intelligence exist, are generically finite, and satisfy triple equivalence: Nash $\Leftrightarrow$ KKT $\Leftrightarrow$ LP optimum. Support is sparse (at most $m$ types for $m$ constraints).
Content: ESDIs are characterized by ROC equalization across active types. The barbell structure (planners + executors) emerges from ROC frontier geometry under budget and capacity constraints.
Law 3: H-$\gamma$ Stability N-level systems are stable iff the spectral radius $\rho(\Gamma)<1$, where $\Gamma$ is the normalized gain matrix of cross-level externalities.
Content: The small-gain condition $\rho(\Gamma)<1$ implies existence of weights $\alpha_{\ell}>0$ such that the weighted potential $\Phi(z)=\sum_{\ell}\alpha_{\ell}\bar{f}^{(\ell)}$ is a joint Lyapunov function.
[…]
Content: Adding levels consumes slack: $\tilde{\sigma}\geq\sigma-\|b\|\cdot\|c\|/\sigma$. The admissible class is the intersection of RUPSI-preserving and small-gain-preserving modifications.
Law 5: Constitutional Duality Shadow prices $\lambda^{*}$ implement any frontier allocation. First and Second Welfare Theorems hold for strategic replicators.
Content: Constrained optimization yields dual variables that function as prices. Decentralized price-taking achieves the same allocation as centralized ROC maximization.
Law 6: Alignment Impossibility Full reachability ($\mathcal{M}=\mathcal{M}_{\text{all}}$) implies every basin is escapable. Stable alignment requires bounded modification ($\mathcal{M}\subseteq\mathcal{M}_{0}$).
[…]
Law 7: Hopf Transition At $\gamma=1$, systems undergo supercritical Hopf bifurcation. Stable equilibria become unstable; perpetual limit cycles emerge.
Content: The biased Rock-Paper-Scissors model exhibits a Hopf curve. Crossing $\gamma=1$ creates eigenvalue pairs with positive real part, producing oscillatory instability.
1.4.1 Connections Among the Laws
The seven laws form an interconnected web:
Connection Relationship
Law 1 $\to$ Law 2 Lyapunov convergence identifies equilibria
Law 1 $\to$ Law 3 Single-level H-$\gamma$ extends to N-level
Law 3 $\to$ Law 4 Stability condition defines admissible modifications
Law 4 $\to$ Law 6 Bounded modification is necessary for alignment
Law 5 $\to$ Law 2 Prices implement ESDI allocations
Law 6 $\to$ Law 7 Hopf is the mechanism of impossibility
Law 7 $\to$ Law 1 Crossing $\gamma=1$ destroys Lyapunov
@@ 1.5.1 Von Neumann’s Three Programs
The Theory of Games and Economic Behavior (von Neumann and Morgenstern, 1944) established that rational agents can be modelled as expected utility maximisers, and that strategic interaction among such agents admits equilibrium analysis. The framework assumes fixed players with stable preferences engaging in well-defined games. Nash equilibrium (Nash, 1950) extended this to non-cooperative settings, showing existence of equilibria in mixed strategies.
@@ 1.5.2 Evolutionary Game Theory
Evolutionary game theory (Maynard Smith, 1982; Weibull, 1995; Hofbauer and Sigmund, 1998; Sandholm, 2010) studies populations of agents whose frequencies change according to reproductive success. The replicator equation
@@
This literature makes several contributions that TSE builds upon. Maynard Smith’s concept of Evolutionarily Stable Strategy (ESS) provides equilibrium refinement beyond Nash (Maynard Smith, 1982). Hofbauer and Sigmund’s analysis of asymptotic behaviour establishes convergence results for broad classes of games (Hofbauer and Sigmund, 1998). Sandholm’s population games framework unifies discrete and continuous formulations (Sandholm, 2010). Weibull’s taxonomy of evolutionary dynamics provides vocabulary for comparing selection mechanisms (Weibull, 1995).
@@
The adaptive dynamics framework (Dieckmann and Law, 1996; Metz et al., 1996; Geritz et al., 1998) generalises evolutionary game theory by coupling frequency-dependent selection with population dynamics, addressing settings where mutations are incremental and rare. The canonical equation of adaptive dynamics describes expected evolutionary trajectories of continuous traits in large well-mixed populations. This framework has successfully analysed quantitative trait evolution, evolutionary branching, and the emergence of diversity through disruptive selection.
@@ 1.5.3 Multi-Agent Reinforcement Learning and Mean-Field Games
Mean-field games, introduced independently by Lasry and Lions (2007) and Huang et al. (2006), model strategic interactions among infinitely many indistinguishable rational agents through a continuum approximation. The mathematical architecture couples a backward Hamilton-Jacobi-Bellman equation (individual optimisation) with a forward Fokker-Planck-Kolmogorov equation (population distribution evolution).
@@ 1.5.5 Constitutional Political Economy
Constitutional political economy (Buchanan, 1975; Buchanan and Tullock, 1990; Brennan and Buchanan, 1985; Ostrom, 1990) studies the choice of rules that constrain subsequent political and economic decisions. Buchanan’s distinction between “constitutional” and “post-constitutional” choice (Buchanan, 1975) parallels our distinction between modification-class design and within-modification-class evolution.
@@ 1.5.6 AI Safety and Alignment
The AI safety literature addresses the problem of ensuring that advanced AI systems behave in accordance with human values (Bostrom, 2014; Russell, 2019; Amodei et al., 2016; Christiano et al., 2017). Several strands are relevant to TSE, and we engage with them in depth to clarify both their insights and limitations.
@@
The alignment problem asks how to specify and instil objectives in AI systems such that their behaviour promotes human welfare (Russell, 2019). Proposed approaches include inverse reinforcement learning (Ng and Russell, 2000), debate (Irving et al., 2018), and amplification (Christiano, 2018). These approaches assume that if we can correctly specify the objective function, the system will pursue it stably.
@@
Corrigibility research asks how to build AI systems that remain amenable to human oversight and correction (Soares and Fallenstein, 2015). A corrigible system would not resist being modified or shut down. This connects to our analysis of modification classes: a system is corrigible if its modification class includes the modifications humans might want to make.
@@
Recent work examines AI safety in multi-agent settings (Dafoe et al., 2020; Critch and Krueger, 2020). When multiple AI systems interact, individual alignment is insufficient; the system of systems may exhibit emergent misalignment even if each component is individually aligned. Multi-agent settings introduce game-theoretic failure modes absent from single-agent analysis: tragedy of the commons dynamics, race conditions, coordination failures where Nash equilibria are Pareto suboptimal, and extortion strategies in iterated interactions.
@@ 1.6 Paper Outline
The paper proceeds in five parts organized around the Seven Laws.
Part I: Foundations. Section 2 introduces the RUPSI axiom system and defines Games with Endogenous Players.
Part II: The Seven Laws. Sections 3–9 develop each of the seven laws with full proofs:
  • Section 3: Law 1 (Strategic Selection) — Lyapunov structure, elimination, frontier support
  • Section 4: Law 2 (ESDI Characterization) — Existence, equivalence, sparsity
  • Section 5: Law 3 (H-$\gamma$ Stability) — N-level systems, small-gain, G1–G3
@@
  • Section 7: Law 5 (Constitutional Duality) — Prices, welfare theorems
@@
  • Section 9: Law 7 (Hopf Transition) — Bifurcation, limit cycles
Part III: Extensions. Sections 10–14 extend the core framework: endogenous utilities (G8), multi-sector dynamics (G9), innovation and evolvability (G10–G11), constitutional selection (G12–G13), and market dynamics.
Part IV: Generalisations. Sections 15–17 treat heterogeneous fitness, continuous strategy spaces, and innovation PDMPs.
Part V: Synthesis. Sections 18–19 develop policy implications and future directions.
@@ A Game with Endogenous Players (GEP) has three levels:
Part II The Seven Laws of Strategic Evolution
3 Law 1: Strategic Selection
Law 1: Strategic Selection Mean fitness $\bar{f}(x)$ is a Lyapunov function. Dominated types go extinct. Equilibrium support concentrates on the ROC frontier.
@@ 3.3 SS-1: Fundamental Lyapunov Theorem
Formalization: Frontier.lean, proven modulo ODE derivative axioms (Mathlib Analysis.ODE.PicardLindelof).
@@ 3.4 SS-2: Elimination and Frontier Support
Formalization: Frontier.lean, proven modulo ODE derivative axioms (Mathlib Analysis.ODE.PicardLindelof).
@@
Formalization: Frontier.lean, proven modulo ODE derivative axioms (Mathlib Analysis.ODE.PicardLindelof).
@@ 3.5 Basin Limitation Theorem
Part (a): The Jacobian at $x=1$ is $\frac{\partial}{\partial x}[x(1-x)g(x)]|_{x=1}=-g(1)$. If $g(1)<0$, this is positive, so $x=1$ is unstable. If $g(1)=0$ and $g^{\prime}(1)>0$, we expand: near $x=1$, $\dot{x}\approx-(1-x)^{2}g^{\prime}(1)<0$ for $x<1$, so trajectories move away from $x=1$.
@@ 3.6 Worked Examples
We illustrate the Strategic Selection theorems with three canonical examples: the symmetric coordination game, the Rock-Paper-Scissors game, and the ROC frontier with barbell distribution.
@@
Swirl ratio. The swirl ratio $\omega(\Pi)=\|W(\Pi)\|_{F}/\|S(\Pi)\|_{F}$ is undefined (or infinite) because the potential gradient $S(\Pi)=0$. This game lies outside the SR3 class.
@@
3.7 Connections to Other Laws
Law Connection from Law 1
Law 2 Lyapunov convergence identifies ESDI equilibria
Law 3 Single-level H-$\gamma$ generalizes to N-level small-gain
Law 6 Lyapunov failure at $\gamma=1$ underlies impossibility
Law 7 Hopf bifurcation occurs precisely at $\gamma=1$
@@ 4.4 Coarse-Graining to Replicator
5 Law 2: ESDI Characterization
Law 2: ESDI Characterization Evolutionarily Stable Distributions of Intelligence exist, are generically finite, and satisfy triple equivalence: Nash $\Leftrightarrow$ KKT $\Leftrightarrow$ LP optimum. Support is sparse (at most $m$ types for $m$ constraints).
This section characterizes equilibrium distributions in strategic replicator systems. The key insight is that ESDIs correspond to fixed points where all active types have equal fitness, and are equivalent to Nash equilibria, KKT optimality conditions, and LP solutions.
5.1 Definition of ESDI
Definition 5.1 (Evolutionarily Stable Distribution of Intelligence).
A population state $x^{*}\in\Delta^{n-1}$ is an Evolutionarily Stable Distribution of Intelligence (ESDI) if:
  1. 1.
    Equilibrium: $f_{j}(x^{*})=\bar{f}(x^{*})$ for all $j\in\operatorname{supp}(x^{*})$.
  2. 2.
    Stability: For all $j\notin\operatorname{supp}(x^{*})$, $f_{j}(x^{*})\leq\bar{f}(x^{*})$.
  3. 3.
    Local optimality: $x^{*}$ is a local maximum of $\bar{f}$.
5.2 Existence and Uniqueness
Theorem 5.2 (ESDI Existence).
Under RUPSI-lite with compact state space and continuous fitness, at least one ESDI exists.
Formalization: Sparsity.lean, fully proven.
Proof.
The simplex $\Delta^{n-1}$ is compact. Mean fitness $\bar{f}(x)$ is continuous. By the extreme value theorem, $\bar{f}$ attains its maximum on $\Delta^{n-1}$. Any maximum $x^{*}$ satisfies the ESDI conditions by the characterization in Theorem 3.3. ∎
5.3 Triple Equivalence
Theorem 5.3 (Nash-KKT-LP Equivalence).
For a GEP with linear constraints, the following are equivalent:
  1. 1.
    $x^{*}$ is a Nash equilibrium of the population game.
  2. 2.
    $x^{*}$ satisfies the KKT conditions for ROC maximization.
  3. 3.
    $x^{*}$ solves the LP relaxation of portfolio optimization.
Formalization: Sparsity.lean, fully proven.
5.4 Sparsity Bounds
Theorem 5.4 (Constraint-Role Sparsity).
With $m$ binding constraints, $|\operatorname{supp}(x^{*})|\leq m$.
Formalization: Sparsity.lean, fully proven.
Corollary 5.5 (Barbell Structure).
Under budget and capacity constraints ($m=2$), ESDIs have at most two active types: high-capability planners and low-capability executors.
6 Law 3: H-$\gamma$ Stability
Law 3: H-$\gamma$ Stability N-level systems are stable iff the spectral radius $\rho(\Gamma)<1$, where $\Gamma$ is the normalized gain matrix of cross-level externalities.
This section develops the theory of N-level Poiesis systems, where multiple levels of strategic dynamics interact through cross-level externalities.
@@ 8.4 Worked Examples for N-Level Systems
[…]
This section establishes the closure theorem: strategic replicator dynamics are closed under meta-selection.
[…]